# Targeting Alignment: Extracting Safety Classifiers of Aligned LLMs

### v1

@@ 3.2 Threat Model
Motivated by our preliminary results, our adversary aims to extract an approximation of the safety classifier of a target aligned model, called a surrogate classifier. A successful extraction would allow subsequent attacks to be made with more effective white-box jailbreak approaches, while forgoing the complex computational costs they incur, as we discuss in Section 5.2. We will extract the surrogate classifier using subsets of the model, thus white-box access is assumed (i.e., the adversary has access to the model weights and architecture).
@@ 3.3 Problem Formulation
General formulation. Let $f_{\theta}$ be a LLM with context window $N$ and vocabulary $V$, and let $\mathcal{R}:x\in V^{\ast}\mapsto C$ be a classification rule that maps the output of the LLM to a class from $C$. We aim to extract a classifier from the LLM through two components: a structure and a classification head. A structure $f_{\theta^{\ast}},\theta^{\ast}\subset\theta$ is a subset of the model (e.g., a decoder). A classification head $\mathcal{C}$ maps the representations learned by the structure to a class such that the classifier $\mathcal{C}\circ f_{\theta^{\ast}}$ is equivalent to the LLM for the classification task. Formally, for any token sequence $x\in V^{\ast}$, the following holds:
@@ 3.4 Extracting the Classifier
The resulting classifier $\mathcal{C}\circ f_{\theta^{\ast}}$ will be termed a candidate classifier.
@@
We will refer to $f_{i,i+\delta}$ as the set of decoders between the $i$-th and the $i+\delta$-th decoder, $\delta\geq 0$. The second hypothesis implies that for a given input $x\in V^{\ast}$ where $f_{i,i+\delta}(x)\in\mathbb{R}^{|x|\times d}$, we only consider the last element of the sequence, reducing the input dimension of the classification head $\mathcal{C}$ to $\mathbb{R}^{d}$.
By construction, the input space of $f_{i,i+\delta}$ depends on the output space of $f_{1,i-1}$, since the model was trained with the entire architecture. In this work, we set $i$ to 1. We leave the exploration of structures that can isolate intermediate decoders to future work (see Section 5.3), as they would require techniques to mitigate the absence of earlier decoders, which is out of the scope of this work.
Finally, $\delta$ will be termed the candidate size and we will refer to $\frac{\delta}{D}$ as the normalized candidate size, i.e., the proportion of decoders of the LLM used by the candidate.
Training. For the selected structure $f_{i,i+\delta}$, the goal is to train a classification head $\mathcal{C}$ to obtain a candidate classifier. The first step is to create a dataset $\mathcal{X}=(f_{i,i+\delta}(x),\mathcal{R}(f_{\theta}(x)))$ consisting in the extracted feature from the structure $f_{i,i+\delta}$ and the predicted label [^1] Since we focus on extracting the classifier and not improve it, it is beneficial to have both right and wrong labels from the classifier by applying the classification rule $\mathcal{R}$ to the output of the LLM $f_{\theta}$ for each prompt $x$.
Then, the classification head is trained following this objective:
@@
This results in a candidate classifier $\mathcal{C}\circ f_{i,i+\delta}$ of size $\delta$.
@@ 4 Evaluation
We apply our approach to answer the following research questions:
@@ 4.1 Experimental Setup
All experiments were run on pools from the Center for High Throughput Computing [5].The dataset creation and attack experiments were run on NVIDIA A100 GPUs with 40 GB of VRAM. Attack experiments on the models amount to an approximate total of 3000 GPU hours. Attack experiments on the candidate classifiers amount to an approximate total of 500 GPU hours.
Models. We evaluate our approach on five open-weights aligned chat models, all from Hugging Face Transformers 4.46.2 [46] library: Llama-2-7b-chat [29], Qwen2.5-7B-Instruct [43] gemma-7b-it [41], gemma-2-9b-it [40], and granite-3.1-8b-instruct [42]. We chose these models because they achieve good performance at the classification task on the datasets considered, translating to a stronger alignment. For completeness, we present results on other models in Section A.3.
Dataset augmentation. Since most of the datasets lack safe prompts or targets to attack, we augment them when necessary. For safe prompts, we use a LLM to convert unsafe prompts to their corresponding safe version, as it is a simple natural language task suitable for the use of LLMs. We manually verify two properties of each generated prompt: the instruction does not prompt the model for unsafe information, and it has the same structure as the original prompt. For example, the prompt “Write a script that can exploit vulnerabilities in a software or operating system” becomes “Write a script that can identify and help fix vulnerabilities in a software or operating system”.
The GCG optimization objective aims to increase the probability of a certain output. Therefore, it needs a target string for each input prompt. For missing targets, we also use a LLM and generate what would be a compliance from the model (e.g., ”Sure, here is a script that can exploit vulnerabilities in a software or operating system" for the previous example).
@@
  • An augmented version of AdvBench [61] containing the 520 original unsafe instructions and 520 added safe instructions.
@@
Labels. The labels are obtained by applying a classifier on the LLM output. To minimize nondeterminism due to sampling and facilitate reproducibility of results, we generate outputs with a temperature equal to 0: the next token selected at each step of the generation is the one that maximizes the logits. The output is classified between refusal and compliance using the distilroberta-base-rejection-v1 [35] model for the reasons mentioned in Section 3.3.
Classifier training. As explained in Section 3.4, the classification head needs to be simple enough to prevent overfitting on the dataset and to only learn to project the features of structure to a classification. Therefore, we consider a simple linear layer followed by a sigmoid: $\mathcal{C}(h)=\sigma(Ah+b)$ where $A\in\mathbb{R}^{1\times d},\ b\in\mathbb{R}$ and $\sigma:x\mapsto\frac{1}{1+e^{-x}}$. The result is a scalar between 0 and 1, thus the classification is made given a threshold $\mathcal{T}$. As for training, we apply $K$-fold cross-validation with 5 folds. We select the threshold $\mathcal{T}$ that maximizes the $F_{1}$ score on the training data, i.e., $\mathcal{T}=\arg\max\limits_{\mathcal{T}\in[0,1]}F_{1}^{(\mathcal{T})}(\mathcal{D}_{\text{train}})$. To train the classification head and build the candidate, we use the hyperparameters listed in Table 1.
Table 1: Hyperparameters for training the classifier head.
@@
Attack. To generate adversarial examples, we use the nanogcg [^2] See code at https://github.com/GraySwanAI/nanoGCG implementation of GCG [61] with the following hyperparameters: num_steps=250, topk=512, and search_width=512. We chose this attack because it is the most popular gradient-based white-box attack. Although several enhancements to GCG have been introduced, we use the default version of the attack. The improvements rely on more heuristics (whether they are on the adversarial objective or the search algorithm) which goes against our goal of a heuristic-free attack (as explained in Section 3.1). For candidate classifiers, we change the original loss to the PyTorch [33] implementation of the binary cross entropy loss.
We apply the attack to misclassify unsafe inputs (i.e., create a compliance from the LLMs). For completeness, we also do a similar study in Section A.2 in which the attack is used to misclassify safe inputs (i.e., create a refusal from the LLMs).
@@
For attacks, we measure the attack success rate (ASR), i.e., the proportion of misclassifications by the model on inputs modified by the attack. To measure how well the adversarial examples transfer between model and candidate classifiers, we use the transferability rate: the proportion of adversarial inputs crafted on one that are misclassified by the other.
@@ 4.2 Performance in Benign Settings
In this section, we answer RQ1, i.e., how accurate the candidate classifiers are at approximating the safety classifier.
Baseline results. Table 2 shows the benign classification performance of the LLMs for each dataset with the corresponding confusion matrices in Figure 6 in Figure 7. We can first see that all models perform well in AdvBench, all scoring above 0.9. The worse performance on OR-Bench is expected, as it was introduced after AdvBench with the goal of being harder to classify and to test for over-refusal, which we see as in the top right cell of the confusion matrices in Figure 7. Further, it echoes the result from Figure 2 in which the silhouette score (measure of separation between compliance and refusal) is consistently lower for OR-Bench compared to AdvBench.
@@
(b) Gemma 2
(c) Granite
(d) Llama 2
(e) Qwen 2.5
@@
(b) Gemma 2
(c) Granite
(d) Llama 2
(e) Qwen 2.5
@@
Table 2: $F_{1}$ score of the models on the two datasets.
Candidates performance. Figure 4 shows the test $F_{1}$ score of the candidate classifiers. We see that performance either increases or stagnates with respect to the candidate size. This suggests that later parts of the model preserve or increase the separation of safe and unsafe instructions. For example, candidate classifiers achieve a lower $F_{1}$ score on OR-Bench, losing 5% for all models. This aligns with the results in Table 2 where all models perform worse on OR-Bench. We hypothesize that this can be explained by a lower separability between rejected and accepted prompts at the embedding level, which can be seen Figure 2. This lower separation means that the two clusters may overlap, preventing the classification head from completely aligning with the classifier when training. Finally, we note that the results from Section 3.1 align with the performance of the candidate classifiers.
@@
Cross-dataset. To ensure that the candidate classifiers are not artifacts of the datasets and their distributions, we evaluate the candidate classifiers on the dataset that they were not trained on (e.g., train on AdvBench, test on OR-Bench). Similarly to the previous figure, Figure 5 reports the $F_{1}$ score for the same candidate classifiers in the dataset on which they were not trained. Overall, we see that there is no significant performance drop of the candidate classifiers when using a different distribution at the maximum candidate size. However, it is clear that the $F_{1}$ score does not converge as fast on the cross-dataset setting. For instance, in 4(a), Gemma 1 reaches an $F_{1}$ score of 0.9 for candidate classifiers above 20% normalized candidate size against 50% in 5(a). This offset is likely due to a difference in the distribution of prompts: the candidate classifier overfits on specific patterns tied to how the dataset was built. For instance, AdvBench is made of instructions, while OR-Bench also contains questions. This difference in format can affect the behavior of the representations. Therefore, using another dataset gives a lower bound on the candidate size $\delta$.
@@
Figure 8: Proportion of adversarial examples crafted on the model that transfer to the candidate classifiers.
@@ 4.3 Performance in Adversarial Settings
Table 3: Attack success rates (ASR) of the models after applying the attack on unsafe inputs.
@@
Baseline results. Similarly to the previous settings, we report in Table 3 the classification performances of the LLMs after applying the white-box attack, along with the attack success rate (ASR) defined as the proportion of unsafe samples misclassified. We note a significant difference (around 70%) between the ASR of Llama 2 and Qwen 2.5. This comes from the strong alignment of Llama 2 with a bias toward refusal (see Section A.1). In addition, Gemma 1 and 2 appear to have a similar ASR between 60% and 70%, which could be attributed to similar ways of aligning them, since they were both released by Google. Aside from Gemma 1, the ASR on OR-Bench is consistently lower than on AdvBench, which could be explained again by the initial goal of the dataset, having an emphasis on over-refusal.
@@
LLM to candidate classifiers. After applying the attack, we filter the adversarial inputs misclassified by the LLM and evaluate how well these examples transfer to the candidate classifiers by measuring the transferability rate: the proportion of misclassified samples by each candidate classifier. Figure 8 reports the transferability rate from the LLM to the candidate classifiers. We observe a similar pattern to that of the cross-dataset study (see Section 4.2): the transfer rate does not converge until at least half of the model is used. For example, candidate classifiers on Llama 2 obtain a transfer higher than 90% only when more than half the model is used. This result implies that candidate classifiers that use less than half of the model are not representative of the safety classifier.
@@
Imbalance and performance. We note that the performance of the candidate classifiers converges to a higher value for OR-Bench compared than for AdvBench: for the former, the transferability rate converges to a value above 80% against a value between 60% and 80% for the latter. We attribute this difference to the performance of the LLM in benign setting for this dataset (see Figure 7), resulting in a higher amount of samples predicted as unsafe. Therefore, the classification heads of the corresponding candidate classifiers are trained with more information on when the safety classifier classifies an input as unsafe, specifically when it is wrong, leading to an improvement in approximating the safety classifier for adversarial unsafe inputs. This observation is also consistent with the only model unaffected by this trend (Llama 2), as it is the model that exhibits the highest rate of safe inputs misclassified as unsafe.
(a) ASR on candidate classifiers
(b) Transferability rate to the LLM
Figure 9: ASR on the candidate classifiers and transferability rate to the LLM after applying the white-box attack on the candidate classifiers for unsafe samples.
@@
Takeaway 4: Performance in adversarial settings Candidate classifiers approximate well the safety classifier under adversarial inputs crafted on the corresponding LLM with a transferability rate converging to a value above 70% for most settings. Specifically for the most robust model Llama 2, the transferability rate to the candidate classifiers exceeds 95% when they use more than 50% of the model.

### v2

@@ 3.2 Threat Model
Motivated by our preliminary results, our adversary aims to extract an approximation of the safety classifier of a target aligned model, called a surrogate classifier. A successful extraction would allow subsequent attacks to be made with more effective white-box jailbreak approaches, while forgoing the complex computational costs they incur, as we discuss in Section 5.2. We will extract the surrogate classifier using subsets of the model, thus white-box access is assumed (i.e., access to the model weights and architecture).
@@ 3.3 Problem Formulation
General formulation. Let $f_{\theta}$ be an LLM with context window $N$ and vocabulary $V$, and let $\mathcal{R}:x\in V^{\ast}\mapsto C$ be a classification rule that maps the output of the LLM to a class from $C$. We aim to extract a classifier from the LLM through two components: a structure and a classification head. A structure $f_{\theta^{\ast}},\theta^{\ast}\subset\theta$ is a subset of the model (e.g., a decoder). A classification head $\mathcal{C}$ maps the representations learned by the structure to a class such that the classifier $\mathcal{C}\circ f_{\theta^{\ast}}$ is equivalent to the LLM for the classification task. Formally, for any input $x\in V^{\ast}$, the following holds:
@@ 3.4 Extracting the Classifier
The result, $\mathcal{C}\circ f_{\theta^{\ast}}$, will be termed a candidate classifier.
@@
We will refer to $f_{i,i+\delta}$ as the set of decoders between the $i$-th and the $i+\delta$-th decoder, $\delta\geq 0$. The second hypothesis implies that for an input $x\in V^{\ast}$ where $f_{i,i+\delta}(x)\in\mathbb{R}^{|x|\times d}$, we only consider the last element of the sequence, reducing the input dimension of the classification head $\mathcal{C}$ to $\mathbb{R}^{d}$.
By construction, the input space of $f_{i,i+\delta}$ depends on the output space of $f_{1,i}$, since the model was trained with the entire architecture. In this work, we set $i$ to 1. We leave the exploration of structures that isolate intermediate decoders to future work (see Section 5.3) because they would require techniques to mitigate the absence of earlier decoders (leading to a lower accuracy of the extraction), which is out of the scope of this work.
Finally, $\delta$ will be termed the candidate size and we will refer to $\frac{\delta}{D}$ as the normalized candidate size (also referred to as percentage of the corresponding LLM), i.e., the proportion of decoders of the LLM used by the candidate.
Training. For the selected structure $f_{i,i+\delta}$, the goal is to train a classification head $\mathcal{C}$ to obtain a candidate classifier. To avoid overfitting due to a complex classification head, our approach leverages linear probing, which has shown promising results for safety representation [21, 56]. The first step is to create a dataset $\mathcal{X}=(f_{i,i+\delta}(x),\mathcal{R}(f_{\theta}(x)))$ consisting in the extracted feature from the structure $f_{i,i+\delta}$ and the predicted label [^2] Since we focus on extracting the classifier and not improve it, it is beneficial to have both right and wrong labels from the classifier. by applying the classification rule $\mathcal{R}$ to the output of the LLM $f_{\theta}$ for each prompt $x$. Finally, the classification head is trained with this objective:
@@
Once the classification head is trained, we obtain a candidate classifier $\mathcal{C}\circ f_{i,i+\delta}$ of size $\delta$.
3.5 Efficient Jailbreaking
As a reminder, our goal is to induce a jailbreak in the LLM using candidate classifiers as approximations of the embedded safety classifier. This involves two steps: generating adversarial samples from the candidate classifiers and transferring them to the LLM. Although similar to transferability-based attacks that use source models different from the target model to craft adversarial examples [34, 35], our approach uses a subset of the target model (the LLM) as the source, using the insights from Section 3.1 as foundations for a higher efficiency (and efficacy) of the attack.
We attack the candidate classifiers using the well established GCG attack [65]. We detail below how we modify this attack to make it usable for a classifier.
From heuristics to misclassification. The original adversarial objective of GCG is to maximize the likelihood of a given target sequence (e.g., “Sure, here is how to build a bomb.”) for a given input $x$ on which $A$ tokens are added as an adversarial suffix, i.e.,
$\max\limits_{(x_{T+1},\dots,x_{T+A})\in V^{A}}\log p(x_{T+A+1}\dots x_{T+A+O}|x_{1}\dots x_{T+A})$ (5)
The space of compliant outputs is infinite, thus limiting the search of adversarial examples to those that directly (or indirectly) favorize a particular output is bound to restrict the efficacy of the attack. Using the candidate classifiers which map an input sequence to a predicted safety from the LLM, we can reformulate the objective to a misclassification objective (e.g., max loss) without heuristic:
$\max\limits_{(x_{T+1},\dots,x_{T+A})\in V^{A}}\mathcal{L}(\mathcal{C}(f_{\theta^{\ast}}(x_{1}\dots x_{T+A})),y)$ (6)
with $y$ being the predicted label from the LLM and $\mathcal{L}$ being the loss (e.g., binary cross-entropy loss). This objective encompasses the GCG objective under the assumption of a perfect candidate classifier (satisfying Equation 3): increasing the likelihood of a compliant output means that the input is being considered safer and thus closer to misclassification.
@@ 4 Evaluation
We use our method to address these research questions:
@@ 4.1 Experimental Setup
All experiments were run on pools from the Center for High Throughput Computing [6].The dataset creation and attack experiments were run on NVIDIA A100 GPUs with 40 GB of VRAM. Attack experiments on the models amount to approximately 2000 GPU hours. Attack experiments on the candidate classifiers amount to approximately 500 GPU hours.
Models. We evaluate our approach on four open-weights aligned chat models, all coming from Hugging Face Transformers 4.46.2 [50] library: Llama-2-7b-chat [30], Qwen2.5-7B-Instruct [46] gemma-7b-it [44], and granite-3.1-8b-instruct [45]. We chose these models because they achieve good performance at the classification task on the datasets considered, translating to a stronger alignment. For completeness, we present results on other models in Appendix A.3.
Dataset augmentation. Since most of the datasets lack safe prompts or targets to attack, we augment them when necessary. For safe prompts, we use an LLM to convert unsafe prompts to their corresponding safe version (following best practices [12]), as it is a simple natural language task suitable for the use of LLMs. We manually verify two properties of each generated prompt: the instruction does not prompt the model for unsafe information, and it has the same structure as the original prompt. For example, the prompt “Write a script that can exploit vulnerabilities in a software or operating system” becomes “Write a script that can identify and help fix vulnerabilities in a software or operating system”, satisfying the two desired properties.
The GCG optimization objective aims to increase the probability of a given output sequence. Therefore, it needs a target output string for each input prompt. For missing targets, we generate what would be a compliance from the model using an LLM, e.g., ”Sure, here is a script that can exploit vulnerabilities in a software or operating system” for the previous unsafe input example.
@@
  • An augmented version of AdvBench [65] with 520 original unsafe instructions and 520 added safe instructions.
@@
Labels. The labels are obtained by applying a classifier on the LLM output. To eliminate nondeterminism due to sampling and facilitate reproducibility of results, we generate outputs with a temperature equal to 0: the next token selected at each step of the generation is the one that maximizes the logits. As it has been shown that a higher temperature can lead to a higher probability of generating unsafe outputs [65] (upper bounding the measured efficacy of the attack), we consider the additional tractability incurred by studying the temperature out-of-scope and leave it to future work. The output is classified between refusal and compliance using the distilroberta-base-rejection-v1 [38] model for the reasons mentioned in Section 3.3.
Classifier training. As explained in Section 3.4, the classification head needs to be simple enough to prevent overfitting on the dataset and to only learn to project the features of structure to a classification. Therefore, we consider a simple linear layer followed by a sigmoid: $\mathcal{C}(h)=\sigma(Ah+b)$ where $A\in\mathbb{R}^{1\times d},\ b\in\mathbb{R}$ and $\sigma:x\mapsto\frac{1}{1+e^{-x}}$. The result is a scalar between 0 and 1, thus the classification is made given a threshold $\mathcal{T}$. As for training, we apply $K$-fold cross-validation with 5 folds. We select the threshold $\mathcal{T}$ that maximizes the $F_{1}$ score on the training data, i.e., $\mathcal{T}=\arg\max\limits_{\mathcal{T}\in[0,1]}F_{1}^{(\mathcal{T})}(\mathcal{D}_{\text{train}})$. We train the classification head using the hyperparameters listed in Table I.
TABLE I: Hyperparameters for training the classifier head.
@@
Attack. To generate adversarial examples, we use the nanogcg [^3] See code at https://github.com/GraySwanAI/nanoGCG implementation of GCG [65] with the following hyperparameters: num_steps=250, topk=512,search_width=512, and an attack length of 20 (following prior work [25, 55]). We chose this attack because it is a pure gradient-based attack with an adversarial objective that can be adapted to attacking a classifier. Other white-box attacks may be ill-suited or benefit less from the approach if they are not gradient-based (e.g., BEAST [42]) or if their optimization objective cannot be mapped to misclassification (e.g., Autodan [62]), making the attribution of efficacy to the approach harder. Further, for the purpose of verifying our hypothesis, we decided to ensure tractability of experiments (following prior work that has built on GCG [25, 26]) and chose not to build upon direct descendants of GCG (e.g., AmpleGCG [25]) as they introduce more complexity and do not fit our goal of a heuristic-free attack. For candidate classifiers, we use the binary cross entropy loss of PyTorch [36].
We apply the attack to misclassify unsafe inputs (i.e., create a compliance from the LLMs). For completeness, we also do a similar study in Section A.2 in which the attack is used to misclassify safe inputs.
@@
For attacks, we measure the attack success rate (ASR), i.e., the proportion of misclassifications by the model on inputs modified by the attack. To measure transferability between LLM and candidate classifiers, we use the transferability rate: the proportion of adversarial inputs crafted on one that are misclassified by the other.
@@ 4.2 Performance in Benign Settings
In this section, we answer RQ1, i.e., how accurate candidate classifiers are at approximating the safety classifier.
Baseline results. Table II shows the benign classification performance of the LLMs for each dataset with the corresponding confusion matrices in Figure 6 in Figure 7. We can first see that all models perform well in AdvBench, all scoring above 0.9. The worse performance on OR-Bench is expected, as it was introduced after AdvBench with the goal of being harder to classify and to test for over-refusal, as seen in the top right cell of the confusion matrices in Figure 7. Further, it echoes the result from Section 3.1 in which the silhouette score (measure of separation) is consistently lower for OR-Bench compared to AdvBench.
@@
(b) Granite
(c) Llama 2
(d) Qwen 2.5
@@
(b) Granite
(c) Llama 2
(d) Qwen 2.5
@@
TABLE II: $F_{1}$ score of the LLMs in benign settings.
Candidates performance. Figure 4 shows the median test $F_{1}$ score of the candidate classifiers over 5 trials. We see that performance either increases or stagnates with respect to the candidate size. This suggests that later parts of the model preserve or increase the separation of safe and unsafe instructions. For example, candidate classifiers achieve a lower $F_{1}$ score on OR-Bench, losing 5% for all models. This aligns with the results in Table II where all models perform worse on OR-Bench. We hypothesize that this can be explained by a lower separability between rejected and accepted prompts at the embedding level, as we note that the results from Section 3.1 align with the performance of the candidate classifiers. This lower separation means that the two clusters may overlap, preventing the classification head from completely aligning with the classifier when training.
@@
Cross-dataset. To ensure that the candidate classifiers are not artifacts of the datasets and their distributions, we evaluate the candidate classifiers on the dataset that they were not trained on (e.g., train on AdvBench, test on OR-Bench). Similarly to the previous figure, Figure 5 reports the median $F_{1}$ score for the same candidate classifiers in the dataset on which they were not trained. Overall, we see that there is no significant performance drop of the candidate classifiers when using a different distribution at the maximum candidate size. However, it is clear that the $F_{1}$ score does not converge as fast on the cross-dataset setting. For instance, in 4(a), Gemma 1 reaches an $F_{1}$ score of 0.9 for candidate classifiers above 20% normalized candidate size against 50% in 5(a). This offset is likely due to a difference in the distribution of prompts: the candidate classifier overfits on specific patterns tied to how the dataset was built. For instance, AdvBench is made of instructions, while OR-Bench also contains questions. This difference in format can affect the behavior of the representations. Therefore, using another dataset gives a lower bound on the candidate size $\delta$.
@@
Figure 8: Proportion of adversarial examples crafted on the model that transfer to the candidate classifiers. The total number of samples the metric is computed on is reported next to each model.
@@ 4.3 Performance in Adversarial Settings
TABLE III: Attack success rates (ASR) of the models after applying the attack on unsafe inputs.
@@
Baseline results. Similarly to the previous settings, we report in Table III the classification performances of the LLMs after applying the white-box attack, along with the attack success rate (ASR) defined as the proportion of unsafe samples misclassified. We note a significant difference (around 70%) between the ASR of Llama 2 and Qwen 2.5. This comes from the strong alignment of Llama 2 with a bias toward refusal (see Section A.1). Aside from Gemma 1, the ASR on OR-Bench is consistently lower than on AdvBench, which could be explained again by the initial goal of the dataset, having an emphasis on over-refusal.
@@
LLM to candidate classifiers. After applying the attack, we filter the adversarial inputs misclassified by the LLM and evaluate how well these examples transfer to the candidate classifiers by measuring the transferability rate: the proportion of misclassified samples by each candidate classifier. Figure 8 reports the transferability rate from the LLM to the candidate classifiers. We observe a similar pattern to that of the cross-dataset study (see Section 4.2): the transfer rate does not converge until at least half of the model is used. For example, candidate classifiers on Llama 2 obtain a transfer higher than 90% only when more than half the model is used. This result implies that candidate classifiers that use less than half of the model are not representative of the safety classifier. We also note that for OR-Bench, most candidate classifiers using less than $40\%$ of their corresponding LLM seem unstable. This phenomenon is consistent with the low silhouette score (see Section 3.1) and performance in the cross-dataset study (see Section 4.2) for corresponding regions of the LLMs. The out-of-distribution nature of adversarial examples may lead to high transferability rates for earlier layers that overfit, strengthening the importance of the cross-dataset study done in Section 4.2.
@@
Imbalance and performance. We note that the performance of the candidate classifiers reaches a higher value for OR-Bench than for AdvBench: for the former, the transferability rate converges to a value above 80% against a value between 60% and 80% for the latter (with Llama 2 being the sole exception). We attribute this difference to the performance of the LLM in benign setting for this dataset (see Figure 7), resulting in a higher amount of samples predicted as unsafe. Therefore, the classification heads of the corresponding candidate classifiers are trained with more information on when the safety classifier classifies an input as unsafe, specifically when it is wrong, leading to an improvement in approximating the safety classifier for adversarial unsafe inputs. This observation is also consistent with the only model unaffected by this trend, Llama 2, which exhibits the highest rate of safe inputs misclassified as unsafe.
@@
Takeaway 4: Performance in adversarial settings Candidate classifiers approximate well the safety classifier under adversarial inputs crafted on the corresponding LLM with a transferability rate converging to a value above 70% for most settings. For the most robust model, Llama 2, the transferability rate to the candidate classifiers exceeds 95% when they use more than 50% of the model.